OpenStudy (anonymous):

solve show all work: Canoeing. During the first part of a canoe trip, Ken covered 60 km at a certain speed. He then traveled 24 km at a speed that was 4 km/h slower. If the total time for the trip was 8 hr, what was the speed on each part of the trip?

OpenStudy (anonymous):

Sove the following equation for r:\[\frac{60}{r}+\frac{24}{(r-4)}= 8\]\[\left\{r\to \frac{5}{2},r\to 12\right\} \]

OpenStudy (anonymous):

Let v be the\[\frac{60}{v}+\frac{24}{v-4}=8\] speed travelled. For the first part of the trip the time taken was 60/v . For the second leg of the trip the time take was 24/(v-4). The sum of these two is the total time of the trip, given as 8 hours. Solve for v. v is the speed of the first part of the trip, and v-4 is the speed of the second part of the trip.

OpenStudy (anonymous):

r=12, r-4=8

OpenStudy (anonymous):

Thank you for the medals.